OpenStudy (anonymous):

aLever with a output arm of 0.8 meter has a mechanical advantage of 6 What is the lenght of the input arm

OpenStudy (anonymous):

0.8meters * 6

OpenStudy (anonymous):

r_1 * F_1 = r_2 * F_2 r_1 * F_1 = 0.8 * (6 * F_1) therefore r_1 = 0.8 * 6